# Mean-field theory of 1+1D $\mathbb{Z}_{2}$ lattice gauge theory with matter

### v1

@@ III.4 Phase diagram
At low filling, see Fig. 6(a), we find a similar structure in our 1D model, with a deconfined symmetry-protected topological phase at small $\lambda$ transitioning to a confined phase at large $\lambda$, for $h=0$. In contrast to the 2D case, the deconfined topological phase is only robust at $h=0$ and gives way for a confined Higgs phase when $h\neq 0$. Furthermore, the confined phase at $h=0$ is associated with a spontaneously broken symmetry (ferromagnetic ordering of $\mathbb{Z}_{2}$ electric field $\hat{\tau}^{x}$). The latter continuously evolves into a paramagnetic state of $\mathbb{Z}_{2}$ electric field when $h\neq 0$ and remains confined. Finally, on the $\lambda=0$ line for $h\neq 0$ a confined meson Luttinger liquid is realized, which transitions to a deconfined parton Luttinger liquid at the special point $\lambda=h=0$.
@@
Finally, for intermediate densities the spontaneously symmetry-broken phase disappears and the phase diagram only contains an extended confined Higgs phase at $h,\lambda\neq 0$, in addition to the deconfined meson (parton) Luttinger liquid at any $h\neq 0$ (at the special point $h=0$) when $\lambda=0$.
@@ IV.1 Derivation of the mean-field theory
^$\hat{\mathcal{H}}_{\rm MF}^{a}=-t\left\langle\hat{\tau}^{z}_{{\langle i,j\rangle}}\right\rangle\sum_{j}\left(\hat{a}^{\dagger}_{j+1}\hat{a}_{j}+\hat{a}^{\dagger}_{j}\hat{a}_{j+1}\right)\\ +\lambda\left\langle\hat{\tau}^{z}_{{\langle i,j\rangle}}\right\rangle\sum_{j}\left(\hat{a}^{\dagger}_{j+1}\hat{a}^{\dagger}_{j}+\hat{a}_{j}\hat{a}_{j+1}\right)\\ -h\left\langle\hat{\tau}^{x}_{{\langle i,j\rangle}}\right\rangle(L+1)+\mu_{a}\sum_{j}\left(\hat{a}^{\dagger}_{j}\hat{a}_{j}-n\right),$ (8)
@@
^$\hat{\mathcal{H}}_{\rm MF}^{\tau}=-t\sum_{j}\left(\left\langle\hat{a}^{\dagger}_{j+1}\hat{a}_{j}\right\rangle+\left\langle\hat{a}^{\dagger}_{j}\hat{a}_{j+1}\right\rangle\right)\hat{\tau}^{z}_{j,j+1}\\ +\lambda\sum_{j}\left(\left\langle\hat{a}^{\dagger}_{j+1}\hat{a}^{\dagger}_{j}\right\rangle+\left\langle\hat{a}_{j}\hat{a}_{j+1}\right\rangle\right)\hat{\tau}_{j,j+1}^{z}\\ -h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}+\mu_{\tau}\sum_{\langle i,j,k\rangle}\left(\hat{\tau}^{x}_{{\langle i,j\rangle}}\hat{\tau}^{x}_{\langle j,k\rangle}-1-2n\right).$ (9)
@@
^$\hat{\mathcal{H}}_{\rm MF}=-g\sum_{{\langle i,j\rangle}}\hat{\tau}_{{\langle i,j\rangle}}^{z}-h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}+\mu_{\tau}\sum_{j}\hat{\tau}^{x}_{{\langle i,j\rangle}}\hat{\tau}^{x}_{\langle j,k\rangle},$ (11)
@@ IV.2.1 The Ising model with transverse and longitudinal fields
This changes when the longitudinal field is nonzero. For the FM Ising interaction $\mu_{\tau}<0$ we always obtain a disordered phase. However for the AFM Ising interaction $\mu_{\tau}>0$ there exists an AFM lobe which remains stable up to $h/\mu_{\tau}<2$ and $g/\mu_{\tau}<1$ [51, 47]. In addition there is an Ising multi-critical point at $h/\mu_{\tau}=2$ and $g/\mu_{\tau}=0$. Both AFM and FM models have a transition point at the transverse field $g/\mu_{\tau}=1$ in the absence of the longitudinal field $h=0$ [53, 47]. Hence this means the transition between the AFM (FM) and disordered states in the effective model depends on the filling of the $\mathbb{Z}_{2}$ LGT system, which we want to describe.
@@ IV.2.3 Mean-field theory entanglement entropy
However, there are no features visible on the $\lambda=0$ lines in the gauge sector. We attribute this to the fact that the gauge-sector mean-field theory in Eq. (11) does not posses the U(1) symmetry in the charges, since the filling is enforced on average only, via the chemical potential $\mu_{\tau}$. i.e., the Lagrangian multiplier. Hence we cannot obtain the quantum criticality on the $\lambda=0$ lines. However, the matter-sector of the mean-field theory, Eq. (8), is critical on the $\lambda=0$ line where it conserves the global U$(1)$ symmetry. While the mean-field ansatz overall captures free partons in the regime $h=0$, it fails to capture in detail the confined meson Luttinger liquid at $h\neq 0,\lambda=0$.
@@ IV.3.1 Mean-field theory Greens function
$\mathcal{G}(x)=\Big{\langle}\frac{1}{4}\Big{(}\prod_{x_{0}\leq\ell<x}\hat{\tau}^{z}_{\ell,\ell+1}\Big{)}\\ \left(1+\hat{\tau}^{x}_{x-1,x}\hat{\tau}^{x}_{x,x+1}\right)\left(1+\hat{\tau}^{x}_{x_{0}-1,x_{0}}\hat{\tau}^{x}_{x_{0},x_{0}+1}\right)\Big{\rangle}.$ (16)
@@
We thus conclude that the mean-field theory captures the main features of the Green’s function behavior for different values of the $\mathbb{Z}_{2}$ electric field for $\lambda\neq 0$. For $h=0$ it also captures the confinement-deconfinement transition to the spontaneously symmetry breaking phases at low and high fillings. However, such transition is also observed for $\lambda=0$ since there is no U(1) symmetry in the mean-field theory in the gauge sector.
@@ IV.4 Summary of the mean-field theory results
The phase diagram of the mean-field model in the gauge sector Eq. (11) thus qualitatively resembles the exact model with two exceptions. Firstly, there is no parton LL or meson LL, since the U(1) symmetry is explicitly broken. Secondly, the transitions to the symmetry-broken states do not exhibit any dependence on the filling.
@@ V.1 Ground state energy comparison
In order to directly compare to the mean-field theory we calculate the ground state energies of the exact $\mathbb{Z}_{2}$ LGT Hamiltonian (1) and the mean-field Hamiltonian (11) for different values of the $\mathbb{Z}_{2}$ electric fields in Fig. 12. We use DMRG [39, 40] where we subtracted the chemical potential contribution from the overall ground state energy, see Appendix D.3. We observe very good qualitative agreement between both Hamiltonians. To be more precise we observe the same change of the typical free-parton parabola for $h=0$, to a deformed concave curve for $h/t=1$. This means that the effective mean-field theory captures the main features of the ground state energy of the exact $\mathbb{Z}_{2}$ LGT.
Figure 12: Comparison of the ground state energy of the $\mathbb{Z}_{2}$ LGT Hamiltonian (1) and the mean-field Hamiltonian (11) for different values of the pairing term $\lambda$ as a function of filling $n$. (a) The ground state energy is symmetric around half-filling where it has a minimum in the deconfined phase $h=0$, which is captured well by the mean-field model. (b) In the confined phase $h/t=1$ the ground state energy rises monotonically for $\lambda=0$ with filling $n$. A minimum reappears at finite filling when $\lambda\neq 0$. Mean-field theory matches the ground state energies well for higher fillings in the confined state. Smaller circles represent the DMRG results of the exact LGT and slightly larger squares represent the mean-field results. We only present $\lambda\geq 0$ as the energy is invariant for ${\lambda\rightarrow-\lambda}$.
@@ V.2 Electric polarization comparison
In order to better understand the mean-field theory we also consider the polarization of the $\mathbb{Z}_{2}$ electric field. We define the polarization in the $x-$direction as
@@
Finite value of the $\mathbb{Z}_{2}$ electric field polarization thus signals the FM phase.
We observe that the onset of finite polarization as a function of filling $n$ changes with $\lambda$ in the case when $h=0$, see Fig. 13(a). For $|\lambda|>0$ the finite polarization persists for higher values of filling $n$ than for the U(1) conserving case when $\lambda=0$, where no spontaneous polarization is found. Comparing this result directly with the mean-field theory in Fig. 13(c) we see that there is no change in polarization as a function of filling $n$ in the latter when we tune the SC term $\lambda$. The drop of polarization to zero occurs at the same value as for $\lambda\neq 0$ in the exact case. This again shows that the mean-field theory correctly captures the qualitative features of the full $\mathbb{Z}_{2}$ LGT, especially when $\lambda\neq 0$, when it starts to match quantitatively.
@@ VI Summary
In this work we develop a mean-field theory for a 1+1D $\mathbb{Z}_{2}$ LGT with a superconducting term which breaks the U(1) symmetry associated with parton number conservation. We first study the phase diagram of the $\mathbb{Z}_{2}$ LGT as a function of the SC term $\lambda$ and filling $n$ at different values of the $\mathbb{Z}_{2}$ electric field $h$, which we sketch in Fig. 1(c) and (d). We use DMRG and analytical mappings to known models, like the Kitaev chain and the transverse field Ising model.
@@
When the $\mathbb{Z}_{2}$ electric field term is non-zero, $h\neq 0$, the partons always confine into mesons in $1+1$D. For $\lambda=0$ mesons form a LL which is protected by the U(1) symmetry. For $\lambda\neq 0$ the system forms a confined Higgs phase, which undergoes a transition to a symmetry-broken AFM state for high filling $n>0.85$.
@@ A.1 Mapping of the $\mathbb{Z}_{2}$ LGT at $h=0$ to the 1D superconducting model
^$\hat{\mathcal{H}}=-t\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}_{j}+\text{h.c.}\right)\\ +\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\hat{n}_{j}.$ (18)
@@
$\hat{b}^{\dagger}_{j}=\Bigg{(}\prod_{l<j}\hat{\tau}^{z}_{l,l+1}\Bigg{)}\hat{a}^{\dagger}_{j},\quad\hat{b}_{j}=\Bigg{(}\prod_{l<j}\hat{\tau}^{z}_{l,l+1}\Bigg{)}\hat{a}_{j},$ (19)
@@
$\hat{a}^{\dagger}_{j}=\Bigg{(}\prod_{l<j}\hat{\tau}^{z}_{l,l+1}\Bigg{)}\hat{b}^{\dagger}_{j},\quad\hat{a}_{j}=\Bigg{(}\prod_{l<j}\hat{\tau}^{z}_{l,l+1}\Bigg{)}\hat{b}_{j}.$ (20)
@@
^$\hat{\mathcal{H}}=-t\sum_{j}\left(\hat{b}^{\dagger}_{j}\hat{b}_{j+1}+\text{h.c.}\right)\\ +\lambda\sum_{j}\left(\hat{b}^{\dagger}_{j+1}\hat{b}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\left(\hat{b}^{\dagger}_{j}\hat{b}_{j}-1/2\right).$ (21)
@@
^$\hat{\mathcal{H}}_{K}=-t\sum_{j}\left(\hat{c}^{\dagger}_{j}\hat{c}_{j+1}+\text{h.c.}\right)\\ -\lambda\sum_{j}\left(\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\left(\hat{c}^{\dagger}_{j}\hat{c}_{j}-1/2\right).$ (24)
@@
^$\hat{\mathcal{H}}=-t\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}_{j}+\text{h.c.}\right)-h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}\\ -\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\hat{n}_{j}.$ (26)

### v2

@@ III.4 Phase diagram
At low filling, see Fig. 6(a), we find a similar structure in our 1D model, with a deconfined symmetry-protected topological phase at small $\lambda$ transitioning to a confined phase at large $\lambda$, for $h=0$. In contrast to the 2D case, the deconfined topological phase is only robust at $h=0$ and gives way for a confined Higgs phase when $h\neq 0$. Furthermore, the confined phase at $h=0$ is associated with a spontaneously broken symmetry (ferromagnetic ordering of $\mathbb{Z}_{2}$ electric field $\hat{\tau}^{x}$). The latter continuously evolves into a paramagnetic state of $\mathbb{Z}_{2}$ electric field term when $h\neq 0$ and remains confined. Finally, on the $\lambda=0$ line for $h\neq 0$ a confined meson Luttinger liquid is realized, which transitions to a deconfined parton Luttinger liquid at the special point $\lambda=h=0$.
@@
Finally, for intermediate densities the spontaneously symmetry-broken phase disappears and the phase diagram only contains an extended confined Higgs phase at $h,\lambda\neq 0$, in addition to the confined meson (parton) Luttinger liquid at any $h\neq 0$ (at the special point $h=0$) when $\lambda=0$.
@@ IV Mean-field theory
We note that variational mean-field approach has been successfully used to study transitions in LGTs in higher dimensions [51, 52, 53].
@@ IV.1 Derivation of the mean-field theory
$\hat{\mathcal{H}}_{\rm MF}^{a}=-t\left\langle\hat{\tau}^{z}_{{\langle i,j\rangle}}\right\rangle\sum_{j}\left(\hat{a}^{\dagger}_{j+1}\hat{a}_{j}+\hat{a}^{\dagger}_{j}\hat{a}_{j+1}\right)\\ +\lambda\left\langle\hat{\tau}^{z}_{{\langle i,j\rangle}}\right\rangle\sum_{j}\left(\hat{a}^{\dagger}_{j+1}\hat{a}^{\dagger}_{j}+\hat{a}_{j}\hat{a}_{j+1}\right)\\ -h\left\langle\hat{\tau}^{x}_{{\langle i,j\rangle}}\right\rangle(L+1)+\mu_{a}\sum_{j}\left(\hat{a}^{\dagger}_{j}\hat{a}_{j}-n\right),$ (8)
@@
$\hat{\mathcal{H}}_{\rm MF}^{\tau}=-t\sum_{j}\left(\left\langle\hat{a}^{\dagger}_{j+1}\hat{a}_{j}\right\rangle+\left\langle\hat{a}^{\dagger}_{j}\hat{a}_{j+1}\right\rangle\right)\hat{\tau}^{z}_{j,j+1}\\ +\lambda\sum_{j}\left(\left\langle\hat{a}^{\dagger}_{j+1}\hat{a}^{\dagger}_{j}\right\rangle+\left\langle\hat{a}_{j}\hat{a}_{j+1}\right\rangle\right)\hat{\tau}_{j,j+1}^{z}\\ -h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}+\mu_{\tau}\sum_{\langle i,j,k\rangle}\left(\hat{\tau}^{x}_{{\langle i,j\rangle}}\hat{\tau}^{x}_{\langle j,k\rangle}-1-2n\right).$ (9)
@@
$\hat{\mathcal{H}}_{\rm MF}=-g\sum_{{\langle i,j\rangle}}\hat{\tau}_{{\langle i,j\rangle}}^{z}-h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}+\mu_{\tau}\sum_{j}\hat{\tau}^{x}_{{\langle i,j\rangle}}\hat{\tau}^{x}_{\langle j,k\rangle},$ (11)
@@ IV.2.1 The Ising model with transverse and longitudinal fields
This changes when the longitudinal field is nonzero. For the FM Ising interaction $\mu_{\tau}<0$ the system remains a FM, and the field lifts the degeneracy between the two possible FM states, according to the sign of $h$ [58]. However for the AFM Ising interaction $\mu_{\tau}>0$ there exists an AFM lobe which remains stable up to $h/\mu_{\tau}<2$ and $g/\mu_{\tau}<1$ [54, 47]. In addition there is an Ising multi-critical point at $h/\mu_{\tau}=2$ and $g/\mu_{\tau}=0$. Both AFM and FM models have a transition point at the transverse field $g/\mu_{\tau}=1$ in the absence of the longitudinal field $h=0$ [56, 47]. Hence this means the transition between the AFM (FM) and disordered states in the effective model depends on the filling of the $\mathbb{Z}_{2}$ LGT system, which we want to describe.
@@ IV.2.3 Mean-field theory entanglement entropy
However, there are no features visible on the $\lambda=0$ lines in the gauge sector. We attribute this to the fact that the gauge-sector mean-field theory in Eq. (11) does not posses the U(1) symmetry in the charges, since the filling is enforced on average only, via the chemical potential $\mu_{\tau}$. i.e., the Lagrangian multiplier. Hence we cannot obtain the quantum criticality on the $\lambda=0$ lines. However, the matter-sector of the mean-field theory, Eq. (8), is critical on the $\lambda=0$ line where it conserves the global U(1)(1) symmetry. While the mean-field ansatz overall captures free partons in the regime $h=0$, it fails to capture in detail the confined meson Luttinger liquid at $h\neq 0,\lambda=0$.
@@ IV.3.1 Mean-field theory Greens function
$\mathcal{G}(x)=\Big\langle\frac{1}{4}\Big(\prod_{x_{0}\leq\ell<x}\hat{\tau}^{z}_{\ell,\ell+1}\Big)\\ \left(1-\hat{\tau}^{x}_{x-1,x}\hat{\tau}^{x}_{x,x+1}\right)\left(1+\hat{\tau}^{x}_{x_{0}-1,x_{0}}\hat{\tau}^{x}_{x_{0},x_{0}+1}\right)\Big\rangle.$ (16)
@@
We thus conclude that the mean-field theory captures the main features of the Green’s function behavior for different values of the $\mathbb{Z}_{2}$ electric field term for $\lambda\neq 0$. For $h=0$ it also captures the confinement-deconfinement transition to the spontaneously symmetry breaking phases at low and high fillings. However, such transition is also observed for $\lambda=0$ since there is no U(1) symmetry in the mean-field theory in the gauge sector.
@@ IV.4 Summary of the mean-field theory results
The phase diagram of the mean-field model in the gauge sector Eq. (11) thus qualitatively resembles the exact model with two exceptions. Firstly, there is no parton LL or meson LL, since the U(1) symmetry is explicitly broken. Secondly, the transitions to the symmetry-broken states do not exhibit any dependence on the filling. This is due to the fact that in the mean-field theory the strength of the SC term $\lambda$, only changes the value of $g$ in the transverse field term $\propto-g\sum_{\langle i,j\rangle}\hat{\tau}^{z}_{\langle i,j\rangle}$. Relating the mean-field theory to the exact mapping of the $\mathbb{Z}_{2}$ LGT to the spin-1/2 model, this means that the mean-field theory effectively remains in the regime where $\lambda=-t$; see Appendix B, for the mapping of the $\mathbb{Z}_{2}$ LGT to the spin-1/2 model.
Figure 12: Comparison of the ground state energy of the $\mathbb{Z}_{2}$ LGT Hamiltonian (1) and the mean-field Hamiltonian (11) for different values of the pairing term $\lambda$ as a function of filling $n$. (a) The ground state energy is symmetric around half-filling where it has a minimum in the deconfined phase $h=0$, which is captured well by the mean-field model. (b) In the confined phase $h/t=1$ the ground state energy rises monotonically for $\lambda=0$ with filling $n$. A minimum reappears at finite filling when $\lambda\neq 0$. Mean-field theory matches the ground state energies well for higher fillings in the confined state. Smaller circles represent the DMRG results of the exact LGT and slightly larger squares represent the mean-field results. We only present $\lambda\geq 0$ as the energy is invariant for $\lambda\rightarrow-\lambda$}. Both legends apply for the two panels in the figure.
@@ V.1 Ground state energy comparison
In order to directly compare to the mean-field theory we calculate the ground state energies of the exact $\mathbb{Z}_{2}$ LGT Hamiltonian (1) and the mean-field Hamiltonian (11) for different values of the pairing term $\lambda$ and the $\mathbb{Z}_{2}$ electric field term $h$ in Fig. 12. We use DMRG [39, 40] where we subtracted the chemical potential contribution from the overall ground state energy, see Appendix D.3. We observe very good qualitative agreement between both Hamiltonians. To be more precise we observe the same change of the typical free-parton parabola for $h=0$, to a deformed concave curve for $h/t=1$. This means that the effective mean-field theory captures the main features of the ground state energy of the exact $\mathbb{Z}_{2}$ LGT.
@@ V.2 Electric polarization comparison
In order to better understand the mean-field theory we also consider the polarization of the $\mathbb{Z}_{2}$ electric field term. We define the polarization in the $x$-direction as
@@
Finite value of the $\mathbb{Z}_{2}$ electric field polarization thus signals the FM phase. We observe that the onset of finite polarization as a function of filling $n$ changes with $\lambda$ in the case when $h=0$, see Fig. 13(a). For $|\lambda|>0$ the finite polarization persists for higher values of filling $n$ than for the U(1) conserving case when $\lambda=0$, where no spontaneous polarization is found. Comparing this result directly with the mean-field theory in Fig. 13(c) we see that there is no change in polarization as a function of filling $n$ in the latter when we tune the SC term $\lambda$. The drop of polarization to zero occurs at the same value as for $\lambda\neq 0$ in the exact case. This again shows that the mean-field theory correctly captures the qualitative features of the full $\mathbb{Z}_{2}$ LGT, especially when $\lambda\neq 0$.
@@ VI Summary
In this work we develop a mean-field theory for a 1+1D $\mathbb{Z}_{2}$ LGT with a superconducting term which breaks the U(1) symmetry associated with parton number conservation. We first study the phase diagram of the $\mathbb{Z}_{2}$ LGT as a function of the SC term $\lambda$ and filling $n$ at different values of the $\mathbb{Z}_{2}$ electric field term $h$, which we sketch in Fig. 1(c) and (d). We use DMRG and analytical mappings to known models, like the Kitaev chain and the transverse field Ising model.
@@
When the $\mathbb{Z}_{2}$ electric field term is non-zero, $h\neq 0$, the partons always confine into mesons in 1+11+1D. For $\lambda=0$ mesons form a LL which is protected by the U(1) symmetry. For $\lambda\neq 0$ the system forms a confined Higgs phase, which undergoes a transition to a symmetry-broken AFM state for high filling $n>0.85$.
@@ A.1 Mapping of the $\mathbb{Z}_{2}$ LGT at $h=0$ to the 1D superconducting model
$\hat{\mathcal{H}}=-t\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}_{j}+\text{h.c.}\right)\\ +\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\hat{n}_{j}.$ (18)
@@
$\hat{b}^{\dagger}_{j}=\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\hat{a}^{\dagger}_{j},\quad\hat{b}_{j}=\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\hat{a}_{j},$ (19)
@@
$\hat{a}^{\dagger}_{j}=\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\hat{b}^{\dagger}_{j},\quad\hat{a}_{j}=\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\hat{b}_{j}.$ (20)
@@
$\hat{\mathcal{H}}=-t\sum_{j}\left(\hat{b}^{\dagger}_{j}\hat{b}_{j+1}+\text{h.c.}\right)\\ +\lambda\sum_{j}\left(\hat{b}^{\dagger}_{j+1}\hat{b}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\left(\hat{b}^{\dagger}_{j}\hat{b}_{j}-1/2\right).$ (21)
@@
$\hat{\mathcal{H}}_{K}=-t\sum_{j}\left(\hat{c}^{\dagger}_{j}\hat{c}_{j+1}+\text{h.c.}\right)\\ -\lambda\sum_{j}\left(\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\left(\hat{c}^{\dagger}_{j}\hat{c}_{j}-1/2\right).$ (24)
@@
$\hat{\mathcal{H}}=-t\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}_{j}+\text{h.c.}\right)-h\sum_{{\langle i,j\rangle}}\hat{\tau}^{x}_{{\langle i,j\rangle}}\\ -\lambda\sum_{{\langle i,j\rangle}}\left(\hat{a}^{\dagger}_{i}\hat{\tau}^{z}_{{\langle i,j\rangle}}\hat{a}^{\dagger}_{j}+\text{h.c.}\right)+\mu\sum_{j}\hat{n}_{j}.$ (26)